# Can a $\xi R\phi^{2}$ non-minimal coupling contribute to the non-relativistic gravitational potential?

### v1

@@ Abstract
We investigate the contribution of the $\sqrt{-g}\xi R\phi^{2}$ non-minimal interaction for a massive scalar to the gravitational potential, found via the non-relativistic 2-2 scattering with one and two graviton exchanges up to one loop order. Such coupling can be most naturally motivated from the renormalisation of a scalar field theory with quartic self interaction in a curved spacetime. This coupling is qualitatively different from the minimal ones, like $\sim\sqrt{G}h^{\mu\nu}T_{\mu\nu}$, as the vertices corresponding to the former does not explicitly contain any scalar momenta, but instead explicitly contains the momentum of the graviton. We show that the insertion of one/two graviton-two scalar non-minimal vertices in the standard 2-2 scattering Feynman diagrams does not yield any contribution for the potential up to one loop, owing to the aforesaid explicit appearance of graviton’s transfer momentum in the numerator of the Feynman amplitude. In order to complement this effect, the most natural choice seems to consider the three graviton vertex generated by the $\sim\Lambda\sqrt{-g}/G$ term in the action, where $\Lambda$ is the cosmological constant. This vertex does not contain explicitly any graviton momentum. With this, and assuming short scale scattering much small compared to the Hubble horizon, we compute the seagull, the vacuum polarisation and the fish diagrams and obtain the non-relativistic scattering amplitudes. They indeed yield a two body gravitational potential at ${\cal O}(\xi\Lambda G^{2})$. We further point out some qualitative differences of this potential with that of the standard result of $\xi=0$.
@@ 1 Introduction
There has been a lot of effort in the past and present to understand the low energy effects of perturbative quantum gravity, even though it is not renormalisable. The motivation behind this is to see its predictions at first few order of the perturbation theory, with the hope that they will be verifiable in a not too far away future. We also hope that such perturbative framework will be smoothly embedded in a complete theory of quantum gravity one day. In particular, the perturbative quantum gravity might tell us whether at all the concept of graviton as a quantum massless spin-2 field is meaningful. We refer our reader to [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] and references therein for various tree and one loop computations in this framework, chiefly focussing on the leading quantum correction to the gravitational or Newton potential, emerging from the graviton exchange of two non-relativistic particles. See also [17, 18, 19] and references therein for computation of bending of light. We also refer our reader to [20, 21, 22, 23] and references therein for computation of the corrections to Newton potential via effective field theory techniques. Moreover, see [24, 25, 26, 27, 28] for inclusion of a cosmological constant at small scales, where the background metric can still be taken to be $\eta_{\mu\nu}$. Finally, see e.g. [29, 30, 31, 32, 33] and references therein for discussions in curved spacetimes like de Sitter.
@@
In this work we wish to investigate the curvature-scalar non-minimal coupling term, $\xi\sqrt{-g}R\phi^{2}$, 2, in the context of perturbative quantum gravity for a massive scalar field, with $\xi$ being the non-minimal coupling parameter. Such coupling naturally arises in the presence of background curvature while renormalising a scalar field theory with quartic self interaction [34]. What will be the contribution of this term at lowest orders in the non-minimal coupling $\xi$ and at one loop, to the correction of the Newton potential between two massive scalar particles? Note that this coupling is qualitatively different from the minimal ones. This is because for a scalar field theory with quadratic kinetic term, in the matter-graviton interaction there can be no derivative of the graviton (such as $\kappa h^{\mu\nu}T_{\mu\nu}$), whereas for the non-minimal interaction, there is no derivative of the scalar at all (6). Due to the latter, we will see that replacing any standard matter-graviton vertex by an ${\cal O}(\xi)$ one in any 2-2 scattering diagram, yields no contribution for the potential at tree or one loop level, owing to the appearance of graviton’s transfer momentum in the numerator of the Feynman amplitude. As we need to consider the two scalar-one graviton and two scalar-two graviton vertices generated by the $\sqrt{-g}\xi R\phi^{2}$ term, the diagrams will have the same topology as that of [3, 4]. Thus in order to find out whether at all this interaction can make any contribution to the non-relativistic potential, we must check whether there is any possibility to ‘nullify’ the effect of the aforementioned derivative or momentum of the graviton. We will see that this is possible in the Einstein gravity, if we consider a cosmological constant and the corresponding ${\cal O}(\Lambda\kappa)$ three-graviton vertex generated by the $4\Lambda\sqrt{-g}/\kappa^{2}$ term, 2. Unlike the usual three graviton vertex generated by $2\sqrt{-g}R/\kappa^{2}$, there will be no derivative of the graviton here. Accordingly, we will see in 3 that the seagull (2), the vacuum polarisation (3) and the fish diagrams (4) will contribute to the gravitational potential, at ${\cal O}(\xi\Lambda\kappa^{4})$. We will further emphasise qualitative differences of the potential thus found from the standard ones ($\xi=0=\Lambda$). We also recall that for a conformally symmetric scalar field ($m^{2}=0,\ \xi=1/6$), static spherically symmetric black hole solution is known to exist since long [35, 36, 37, 38]. To the best of our knowledge, the $\xi\sqrt{-g}R\phi^{2}$ non-minimal coupling is rather less investigated from the quantum gravity perspective. Due to its aforementioned qualitatively different nature from the standard matter-graviton coupling, it seems to be an important job to probe the effect of the first on quantum gravity predictions.
@@
We also note that in general the presence of a $\Lambda$ prevents us from doing quantum field theory in a flat background, and instead one should do it in a de Sitter background, which reads for example in the static coordinatisation,
$ds^{2}=-fdt^{2}+f^{-1}dr^{2}+r^{2}d\Omega^{2},$ (1)
@@
where $f=1-\Lambda r^{2}/3$. However much inside the Hubble horizon ($r\ll\sqrt{3/\Lambda}$), the spacetime can be split into a Minkowski background plus a perturbation. The Schwarzschild-de Sitter spacetime, describing the gravitational filed of a static mass inside the de Sitter universe, is given by $f(r)=1-2MG/r-\Lambda r^{2}/3$ in 1. The region with smaller $r$ values are dominated by the attraction due to the mass $M$, and the larger ones by the repulsion due to $\Lambda$. The attraction decreases monotonically with increasing $r$ whereas the repulsion increases. Accordingly, there is a radius, $R_{\rm TA,max}=(3MG/\Lambda)^{1/3},$ known as the maximum turn around radius, around which the attraction and repulsion balance each other, e.g. [39]. It is clear that we may do a decomposition of flat spacetime plus perturbation inside this radius, to look into short wavelength phenomena. Such decomposition in the presence of a $\Lambda$ has been done earlier in [24, 25, 26, 27, 28].
@@
Before we go into the computations, we will summarise below the basic technical ingredients we will need. We will work with mostly positive signature of the metric in four spacetime dimensions. Our notation for symmetrisation will be : $X_{(\mu\nu)}=X_{\mu\nu}+X_{\nu\mu}$.
@@ 2 The basic ingredients
$\partial_{\mu}\left(h^{\mu}{}_{\nu}-\frac{1}{2}\delta^{\mu}_{\nu}h\right)=0,$
@@
$\partial^{2}\left(h_{\mu\nu}-\frac{1}{2}h\eta_{\mu\nu}\right)-\frac{2\Lambda}{\kappa}\eta_{\mu\nu}=0,$ (4)
@@
where the first term on the right hand side is the usual homogeneous massless spin-2 plane wave solution which we will use for field quantisation, whereas the second term corresponds to the linearised, classical de Sitter perturbation yielding the background curvature, $R=4\Lambda$. The field quantisation with respect to $h_{\mu\nu}|_{\rm hom}$ yields the graviton propagator. We will restrict ourselves to ${\cal O}(\Lambda)$ only. We now make the following leading expansions for the scalar and the $\Lambda$ part of the action,
$S_{\phi}=-\frac{1}{2}\int d^{4}x\left[(\partial\phi)^{2}+m^{2}\phi^{2}\right]+\frac{\kappa}{2}\int d^{4}xh^{\mu\nu}\left[(\partial_{\mu}\phi)(\partial_{\nu}\phi)-\frac{1}{2}\eta_{\mu\nu}\left((\partial\phi)^{2}+m^{2}\phi^{2}\right)\right]$
@@
$-\frac{\kappa^{2}}{2}\int d^{4}x\left[\left(h^{\mu}{}_{\alpha}h^{\alpha\nu}-\frac{1}{2}hh^{\mu\nu}+\frac{h^{2}}{8}\eta^{\mu\nu}-\frac{1}{4}h_{\alpha\beta}h^{\alpha\beta}\eta^{\mu\nu}\right)(\partial_{\mu}\phi)(\partial_{\nu}\phi)+\frac{m^{2}}{4}\left(\frac{h^{2}}{2}-h_{\mu\nu}h^{\mu\nu}\right)\phi^{2}\right]$
$+\frac{\xi}{2}\int d^{4}x\left[\kappa\phi^{2}\partial^{2}h-\kappa^{2}\phi^{2}\left\{-\frac{1}{4}h\partial^{2}h-\frac{1}{2}(\partial^{\lambda}h^{\mu\alpha})(\partial_{\alpha}h_{\mu\lambda})+h^{\mu\nu}\partial^{2}h_{\mu\nu}+\frac{3}{4}(\partial^{\lambda}h^{\mu\nu})(\partial_{\lambda}h_{\mu\nu})\right\}\right]+\ {\cal O}(\kappa^{3},\xi\kappa^{3}),$
$S_{\Lambda}=-\frac{4\Lambda}{\kappa^{2}}\int d^{d}x\left[1+\frac{\kappa h}{2}+\frac{\kappa^{2}}{4}\left(\frac{h^{2}}{2}-h_{\mu\nu}h^{\mu\nu}\right)+\frac{\kappa^{3}}{2}\left(\frac{1}{3}h_{\mu\nu}h^{\nu\lambda}h_{\lambda}{}^{\mu}-\frac{1}{4}hh_{\mu\nu}h^{\mu\nu}+\frac{h^{3}}{24}\right)\right]+\ {\cal O}(\Lambda\kappa^{2}),$ (6)
where we have used $m^{2}\gg\Lambda$, i.e. the Compton wavelength of the scalar is much small compared to the Hubble horizon, as required. The scalar and the graviton propagators read
$\Delta(k)=-\frac{1}{k^{2}+m^{2}};\qquad\Delta_{\mu\nu\alpha\beta}(k)=-\frac{P_{\mu\nu\alpha\beta}}{k^{2}},\qquad P_{\mu\nu\alpha\beta}=\frac{1}{2}\left(\eta_{\mu\alpha}\eta_{\nu\beta}+\eta_{\mu\beta}\eta_{\nu\alpha}-\eta_{\mu\nu}\eta_{\alpha\beta}\right),$ (7)
where $\mu,\nu$ and $\alpha,\beta$ pairwise correspond to two graviton fields, $h_{\mu\nu}$ and $h_{\alpha\beta}$. Note in 6 that there is a term quadratic in $h_{\mu\nu}$ in $S_{\Lambda}$, which could also make a contribution proportional to $\Lambda h_{\mu\nu}$ in 4. However, this is subleading compared to the $\eta_{\mu\nu}$ term. Note also from 5 that the classical term will make contribution beyond linear order in $\Lambda$ in $S_{\Lambda}$, and hence we will take in our calculations $h_{\mu\nu}$ to be purely quantum in the cubic interaction. This term will generate a three graviton vertex of ${\cal O}(\kappa\Lambda)$ with no explicit appearance of the graviton momentum. The usual three graviton vertex originating from the $\sqrt{-g}R/2\kappa^{2}$ term (with momenta $l$, $q$, lql-q along the three lines) reads, e.g. [4],
@@
$V^{(3)\mu\nu}_{\alpha\beta\gamma\delta}(\kappa;l,q)$ $=-\frac{i\kappa}{2}\left[P_{\alpha\beta\gamma\delta}\left(l^{\mu}l^{\nu}+(l-q)^{\mu}(l-q)^{\nu}+q^{\mu}q^{\nu}-\frac{3}{2}\eta^{\mu\nu}q^{2}\right)\right$. (8)
[…]
$\left.+\left\{\left(l^{2}+(l-q)^{2}\right)\left(I_{\alpha\beta}{}^{\mu\sigma}I_{\gamma\delta\sigma}{}^{\nu}+I_{\gamma\delta}{}^{\mu\sigma}I_{\alpha\beta\sigma}{}^{\nu}-\frac{1}{2}\eta^{\mu\nu}{\cal P}_{\alpha\beta\gamma\delta}\right)-\left(I_{\gamma\delta}{}^{\mu\nu}\eta_{\alpha\beta}l^{2}+I_{\alpha\beta}{}^{\mu\nu}\eta_{\gamma\delta}(l-q)^{2}\right)\right\}\right],\qquad$
with
$I_{\mu\nu\lambda\rho}=\frac{1}{2}\left(\eta_{\mu\lambda}\eta_{\nu\rho}+\eta_{\mu\rho}\eta_{\nu\lambda}\right).$
We will also require the standard integral [40]
$I^{r}_{s}=\int\frac{d^{d}k}{(2\pi)^{d}}\frac{(k^{2})^{r}}{(k^{2}-a^{2})^{s}}=\frac{i(-1)^{r+s}}{(4\pi)^{d/2}}\frac{1}{(a^{2})^{s-r-d/2}}\frac{\Gamma(r+d/2)\Gamma(s-r-d/2)}{\Gamma(d/2)\Gamma(s)}.$ (9)

### v2

@@ Abstract
This letter investigates the contribution of the $\sqrt{-g}\xi R\phi^{2}$ non-minimal interaction to the long range gravitational potential for massive scalar fields, found from the non-relativistic limit of the 2-2 scattering amplitude with graviton exchanges. Such coupling is most naturally motivated from the renormalisation of a scalar field theory with quartic self interaction in a curved spacetime. This is qualitatively different from the minimal ones like $\sqrt{G}h^{\mu\nu}T_{\mu\nu}$, as the vertices corresponding to the former does not explicitly contain any scalar momenta, but instead explicitly contains the momentum carried by graviton line. For the minimal vertex, the long range gravitational potential up to one loop $({\cal O}(G),{\cal O}(G^{2}))$ was obtained earlier from the terms non-analytic in the transfer momentum, $q^{-2},\ q^{-1},\ \ln q^{2}$, yielding potentials respectively like $r^{-1}$, $r^{-2}$, $r^{-3}$. However owing to the aforesaid explicit appearance of graviton’s transfer momentum for the non-minimal vertices, the leading contribution in this case comes at ${\cal O}(\xi G^{2})$, and turns out to be subleading compared to even $r^{-3}$. In order to complement this ‘screening’ effect, we consider the three graviton vertex generated by the $\sim\Lambda\sqrt{-g}/G$ term in the action, where $\Lambda$ is the cosmological constant. This vertex does not contain any graviton momentum explicitly. With this vertex, and assuming short scale scattering much small compared to the Hubble horizon, we compute the seagull, the vacuum polarisation and the fish diagrams and obtain the 2-2 scattering amplitudes. The leading two body gravitational potential at ${\cal O}(\xi\Lambda G^{2})$ behaves like $r^{-1}$, even though it is much subleading compared to Newton’s potential due to the appearance of $\Lambda$. We also discuss the scenario where this potential dominates the aforesaid ${\cal O}(\xi G^{2})$ one.
@@ 1 Introduction
There has been a lot of effort in the past and present to understand the low energy effects of perturbative quantum gravity, even though it is not renormalisable. The goal behind this is to see its predictions at the first few orders of the perturbation theory, with the hope that they will be verifiable in a not too far away future. We also hope that such perturbative framework will be smoothly embedded in a complete theory of quantum gravity one day. In particular, the perturbative quantum gravity might tell us whether at all the concept of graviton as a quantum massless spin-2 field is meaningful. We refer our reader to [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] and references therein for various tree and one loop computations in this framework, chiefly focussing on the leading quantum corrections to the long range gravitational potential, emerging from the graviton exchanges between two non-relativistic particles. See [21, 22, 23] and references therein for computation of bending of light. We refer our reader to e.g. [24, 25, 26, 27] and references therein for the computations on the corrections to the Newton potential via effective field theory techniques. See also [28] for a technique for extracting potential from the four point amplitude in the post Minkowski expansion. Moreover, see [29, 30, 31, 32, 33] for inclusion of a cosmological constant at subhorizon scales, where the background metric can still be taken to be $\eta_{\mu\nu}$. Finally, see e.g. [34, 35, 36, 37, 38] and references therein for discussions in curved spacetimes like the de Sitter. We note that the long range predictions of perturbative quantum gravity can be interpreted as an effective field theory framework making well defined physical predictions, irrespective of its ultraviolet completion. This serves as one of the chief motivations behind engaging into such low energy computations associated with Einstein’s or other alternative gravity theories.
@@
In this work we wish to probe the curvature-scalar non-minimal coupling term, $\xi\sqrt{-g}R\phi^{2}$, 2, in the context of long range gravitational potential for massive scalar fields, where $\xi$ is the non-minimal coupling parameter. Such coupling naturally arises in the presence of background curvature while renormalising a scalar field theory with quartic self-interaction [39, 40]. What will be the contribution of this term at the lowest order in $\xi$, at tree and one loop level to the correction of the Newton potential? This coupling is qualitatively different from the minimal ones. This is because for a scalar field theory with quadratic kinetic term, in the matter-graviton interaction there can be no derivative of the graviton (such as $\kappa h^{\mu\nu}T_{\mu\nu}$), whereas for the non-minimal interaction, there is no derivative of the scalar at all (6). For $\xi=0$, the leading gravitational potential from 2-2 scalar scattering amplitudes up to one loop was computed earlier from the terms non-analytic in the transfer momentum carried by gravitons ($q^{-2},\ q^{-1},\ \ln q^{2}$), yielding results up to ${\cal O}(r^{-3})$, e.g. [4]. For $\xi\neq 0$ however, we will see below that replacement of any minimal matter-graviton vertex by a non-minimal vertex results in the appearance of graviton’s transfer momentum in the numerator of the corresponding scattering amplitudes. Consequently, no long range contribution is found at tree level. The leading such contribution is found at ${\cal O}(\xi G^{2})$ for some one loop diagrams, and they are subleading compared to even $r^{-3}$. For the two scalar-one graviton and two scalar-two graviton vertices generated by the $\sqrt{-g}\xi R\phi^{2}$ term, the diagrams will have the same topology as that of $\xi=0$ [3, 4]. Thus in order to understand whether there exists any stronger contribution, we must check if there is any way to complement this ‘screening’ effect. We will see that this is possible in the Einstein gravity, if we consider a cosmological constant and the corresponding ${\cal O}(\Lambda\kappa)$ three graviton vertex generated by the $4\Lambda\sqrt{-g}/\kappa^{2}$ term, 2. Unlike the usual three graviton vertex generated by $2\sqrt{-g}R/\kappa^{2}$, there will be no derivative of the graviton here. Accordingly, we will see in 3 that the seagull (2), the vacuum polarisation (3) and the fish diagrams (4) will contribute to the gravitational potential, at ${\cal O}(\xi\Lambda G^{2})$, with the leading behaviour of $r^{-1}$. We will further emphasise qualitative differences of the potential thus found with the standard ones ($\xi=0=\Lambda$). We also recall that for a conformally symmetric scalar field ($m^{2}=0,\ \xi=1/6$), static spherically symmetric black hole solutions are known to exist since long [41, 42, 43, 44]. To the best of our knowledge, the $\xi\sqrt{-g}R\phi^{2}$ coupling, even though qualitatively different from standard minimal ones, is relatively less investigated from quantum gravity perspective. We refer our reader for example, to [45, 46, 47, 48] for discussion on renormalisation in curved spacetime, asymptotic safety and quantum gravity, and Higgs inflation in the presence of such coupling. Due to its aforementioned qualitatively different nature from the standard matter-graviton coupling, it seems to be an important job to further probe the effect of the first on quantum gravity predictions.
@@
Now, in general the presence of a $\Lambda$ prevents us from doing quantum field theory in a flat background, and instead one should do it in a de Sitter background, which reads, for example, in the static coordinatisation,
$ds^{2}=-fdt^{2}+f^{-1}dr^{2}+r^{2}d\Omega^{2}$ (1)
@@
where $f=1-\Lambda r^{2}/3$. However much inside the Hubble horizon ($r\ll\sqrt{3/\Lambda}$), the spacetime can be split into a Minkowski background plus a perturbation. The Schwarzschild-de Sitter spacetime, describing the gravitational filed of a static mass inside the de Sitter universe, is given by $f(r)=1-2MG/r-\Lambda r^{2}/3$ in 1. The region with smaller $r$ values are dominated by the attraction due to the mass $M$, and the larger ones by the repulsion due to $\Lambda$. The attraction decreases monotonically with increasing $r$, whereas the repulsion increases. Accordingly, there is a scale, $R_{\rm TA,max}=(3MG/\Lambda)^{1/3},$ known as the maximum turn around radius, around which the attraction and repulsion balance each other, e.g. [49]. It is clear that we may do a decomposition of flat spacetime plus perturbation inside this radius, essentially to look into short wavelength phenomena. Such decomposition in the presence of a $\Lambda$ has been done earlier in [29, 30, 31, 32, 33]. One may reasonably anticipate that these short scale effects will be tiny while compared to the $\Lambda=0$ cases hitherto investigated. Nevertheless, we believe probing such effects are important, in order to further understand the nature of the matter-gravity interaction. We also note that in the primordial inflationary scenario and in the infrared regime, one can encounter non-perturbative secular large logarithms while computing the quantum corrections to the two body gravitational potential, e.g. [36] and references therein. Understanding the role of the $\xi\sqrt{-g}R\phi^{2}$ coupling in the inflationary and large scale context can be an important task, which we reserve for the future.
We wish to emphasise that the phrases short scale/short wavelength appearing anywhere in this letter is to be understood as the subhorizon/subHubble scale. Even though the potential is long range, it is valid only up to the scale of the maximum turn around radius of a structure. Beyond this scale the expansion of the universe dominates, and the Hubble flow will essentially invalidate the notion of a potential as an explicit time independent non-relativistic quantity.
In this regard, we also note that while the maximum turn around radius mentioned above can be thought of as a local test of cosmological constant/dark energy [49], the other effects of $\Lambda$ on the gravitational physics inside the maximum turn around radius are expected to be small compared to that of the mass of the gravitating object. $\Lambda$ dominates at very large scales beyond the maximum turn around radius, and contributes to phenomena like the gravitational redshift, or correlations originating from the very early universe, etc.. Nevertheless, with increasing measurement facilities such as the event horizon telescope, one can hope to probe the effect of $\Lambda$ in short scale physics (like the solar system) as well, in the not too far away future, like the perihelion precession, bending of light and the Shapiro delay. This might give us new insight about the nature of the dark energy. Thus attempts to understand such phenomena’s perturbative quantum gravity corrections are also well motivated.
@@
Before we go into the computations, we will summarise below the basic technical ingredients we will need. We will work with the mostly positive signature of the metric in four spacetime dimensions. Our notation for symmetrisation will be : $X_{(\mu\nu)}=X_{\mu\nu}+X_{\nu\mu}$.
@@ 2 The basic ingredients
$\partial_{\mu}\left(h^{\mu}{}_{\nu}-\frac{1}{2}\delta^{\mu}_{\nu}h\right)=0$
@@
$\partial^{2}\left(h_{\mu\nu}-\frac{1}{2}h\eta_{\mu\nu}\right)-\frac{2\Lambda}{\kappa}\eta_{\mu\nu}=0$ (4)
@@
where the first term on the right-hand side is the usual homogeneous massless spin-2 plane wave solution, which will give the field quantisation, whereas the second term corresponds to the linearised, classical de Sitter perturbation yielding the background curvature, $R=4\Lambda$. The quantisation for $h_{\mu\nu}|_{\rm hom}$ yields the graviton propagator. The ground state for the theory will be taken to be the standard Poincarré invariant Minkowski vacuum, subject to the flat background, 3. Note that in the presence of $\Lambda$, this cannot be possible at large scales, due to particle creation and the subsequent instability of the initial vacuum state. This happens if we are concerned with length scales comparable to or higher than the Hubble horizon. Such large scale, infrared effects are essentially non-perturbative and non-equilibrium, so that the standard in-out $S$-matrix formalism of quantum field theory which we will be using here, fails. Scattering has no well defined meaning in such dynamical backgrounds and one needs to do the in-in formalism to compute expectation values or correlation functions, e.g. [38] and references therein. This problem is not relevant to our present scenario, for as we have stated in the preceding section, we are concerned here only up to the length scale of a structure (like a galaxy) defined by its maximum turn around radius. This length scale is much subhorizon and hence we may safely ignore the aforementioned large scale IR effects for our present purpose. This means that the $\Lambda$ term appearing in 5 is much small compared to unity. We will assert this scenario in our calculations in the following by taking the Compton wavelength of the scattering particles to be much small compared to the horizon length. Also, since the current observed value of $\Lambda$ is tiny, ${\cal O}(10^{-52}\ {\rm m}^{-2})$, we will restrict our computations to ${\cal O}(\Lambda)$ only.
We now make the following leading expansions for the scalar and the $\Lambda$ part of the action,
@@
$\begin{split}S_{\phi}=&-\frac{1}{2}\int d^{4}x\left[(\partial\phi)^{2}+m^{2}\phi^{2}\right]+\frac{\kappa}{2}\int d^{4}xh^{\mu\nu}\left[(\partial_{\mu}\phi)(\partial_{\nu}\phi)-\frac{1}{2}\eta_{\mu\nu}\left((\partial\phi)^{2}+m^{2}\phi^{2}\right)\right]\\ &-\frac{\kappa^{2}}{2}\int d^{4}x\left[\left(h^{\mu}{}_{\alpha}h^{\alpha\nu}-\frac{1}{2}hh^{\mu\nu}+\frac{h^{2}}{8}\eta^{\mu\nu}-\frac{1}{4}h_{\alpha\beta}h^{\alpha\beta}\eta^{\mu\nu}\right)(\partial_{\mu}\phi)(\partial_{\nu}\phi)+\frac{m^{2}}{4}\left(\frac{h^{2}}{2}-h_{\mu\nu}h^{\mu\nu}\right)\phi^{2}\right]\\ &+\frac{\xi}{2}\int d^{4}x\left[\kappa\phi^{2}\partial^{2}h-\kappa^{2}\phi^{2}\left\{-\frac{1}{4}h\partial^{2}h-\frac{1}{2}(\partial^{\lambda}h^{\mu\alpha})(\partial_{\alpha}h_{\mu\lambda})+h^{\mu\nu}\partial^{2}h_{\mu\nu}+\frac{3}{4}(\partial^{\lambda}h^{\mu\nu})(\partial_{\lambda}h_{\mu\nu})\right\}\right]+\ {\cal O}(\kappa^{3},\xi\kappa^{3})\\ S_{\Lambda}=&-\frac{4\Lambda}{\kappa^{2}}\int d^{d}x\left[1+\frac{\kappa h}{2}+\frac{\kappa^{2}}{4}\left(\frac{h^{2}}{2}-h_{\mu\nu}h^{\mu\nu}\right)+\frac{\kappa^{3}}{2}\left(\frac{1}{3}h_{\mu\nu}h^{\nu\lambda}h_{\lambda}{}^{\mu}-\frac{1}{4}hh_{\mu\nu}h^{\mu\nu}+\frac{h^{3}}{24}\right)\right]+\ {\cal O}(\Lambda\kappa^{2})\end{split}$ (6)
where we have used $m^{2}\gg\Lambda$, i.e., the Compton wavelength of the scalar is much small compared to the Hubble horizon, as stated above. Note in 6 that there is a term quadratic in $h_{\mu\nu}$ in $S_{\Lambda}$, which could also make a contribution proportional to $\Lambda h_{\mu\nu}$ in 4. However, this is subleading compared to the $\eta_{\mu\nu}$ term, and hence has been ignored. Also as 5 suggests, both quantum $h_{\mu\nu}|_{\rm hom}$ and the classical $\Lambda r^{2}$ terms make contributions in the above. However, we will only be concerned about the quantum effects generated by $h_{\mu\nu}|_{\rm hom}$, as the classical terms cannot contribute in forming a vertex or in carrying momenta. Moreover in $S_{\Lambda}$, the classical $\Lambda r^{2}$ term will make contributions beyond linear order in $\Lambda$, hence are further subleading in our subhorizon scenario. Thus we have a three graviton vertex of ${\cal O}(\kappa\Lambda)$ from $S_{\Lambda}$. There is no explicit appearance of the graviton momentum here, for it contains no derivative of the graviton. This vertex will serve the key role in our following computations.
The scalar and the graviton propagators read
$\Delta(k)=-\frac{1}{k^{2}+m^{2}};\qquad\Delta_{\mu\nu\alpha\beta}(k)=-\frac{P_{\mu\nu\alpha\beta}}{k^{2}};\qquad P_{\mu\nu\alpha\beta}=\frac{1}{2}\left(\eta_{\mu\alpha}\eta_{\nu\beta}+\eta_{\mu\beta}\eta_{\nu\alpha}-\eta_{\mu\nu}\eta_{\alpha\beta}\right)$ (7)
where $\mu,\nu$ and $\alpha,\beta$ pairwise correspond to two graviton fields, $h_{\mu\nu}$ and $h_{\alpha\beta}$.
The usual three graviton vertex originating from the $\sqrt{-g}R/2\kappa^{2}$ term (with momenta $l$, $q$, lql-q along the three lines) reads, e.g. [4],
@@
$\begin{split}V^{(3)\mu\nu}_{\alpha\beta\gamma\delta}(\kappa;l,q)=&-\frac{i\kappa}{2}\left[P_{\alpha\beta\gamma\delta}\left(l^{\mu}l^{\nu}+(l-q)^{\mu}(l-q)^{\nu}+q^{\mu}q^{\nu}-\frac{3}{2}\eta^{\mu\nu}q^{2}\right)\right.\\ &\left.+2q_{\lambda}q_{\sigma}\left(I_{\alpha\beta}{}^{\sigma\lambda}I_{\gamma\delta}{}^{\mu\nu}+I_{\gamma\delta}{}^{\sigma\lambda}I_{\alpha\beta}{}^{\mu\nu}-I_{\alpha\beta}{}^{\mu\sigma}I_{\gamma\delta}{}^{\nu\lambda}-I_{\gamma\delta}{}^{\mu\sigma}I_{\alpha\beta}{}^{\nu\lambda}\right)\right.\\ &\left.+\left\{q_{\lambda}q^{\mu}\left(\eta_{\alpha\beta}I_{\gamma\delta}{}^{\nu\lambda}+\eta_{\gamma\delta}I_{\alpha\beta}{}^{\nu\lambda}\right)+q_{\lambda}q^{\nu}\left(\eta_{\alpha\beta}I_{\gamma\delta}{}^{\mu\lambda}+\eta_{\gamma\delta}I_{\alpha\beta}{}^{\mu\lambda}\right)-q^{2}\left(\eta_{\alpha\beta}I_{\gamma\delta}{}^{\mu\nu}+\eta_{\gamma\delta}I_{\alpha\beta}{}^{\mu\nu}\right)\right.\right.\\ &\left.\left.-\eta^{\mu\nu}q_{\sigma}q_{\lambda}\left(\eta_{\alpha\beta}I_{\gamma\delta}{}^{\sigma\lambda}+\eta_{\gamma\delta}I_{\alpha\beta}{}^{\sigma\lambda}\right)\right\}\right.\\ &\left.+\left\{-2q_{\lambda}\left(I_{\alpha\beta}{}^{\lambda\sigma}I_{\gamma\delta\sigma}{}^{\nu}(l-q)^{\mu}+I_{\alpha\beta}{}^{\lambda\sigma}I_{\gamma\delta\sigma}{}^{\mu}(l-q)^{\nu}+I_{\gamma\delta}{}^{\lambda\sigma}I_{\alpha\beta\sigma}{}^{\nu}l^{\mu}+I_{\gamma\delta}{}^{\lambda\sigma}I_{\alpha\beta\sigma}{}^{\mu}l^{\nu}\right)\right.\right.\\ &\left.\left.+q^{2}\left(I_{\alpha\beta\sigma}{}^{\mu}I_{\gamma\delta}{}^{\nu\sigma}+I_{\gamma\delta\sigma}{}^{\mu}I_{\alpha\beta}{}^{\nu\sigma}\right)+\eta^{\mu\nu}q_{\sigma}q_{\lambda}\left(I_{\alpha\beta}{}^{\lambda\rho}I_{\gamma\delta\rho}{}^{\sigma}+I_{\gamma\delta}{}^{\lambda\rho}I_{\alpha\beta\rho}{}^{\sigma}\right)\right\}\right.\\ &\left.+\left\{\left(l^{2}+(l-q)^{2}\right)\left(I_{\alpha\beta}{}^{\mu\sigma}I_{\gamma\delta\sigma}{}^{\nu}+I_{\gamma\delta}{}^{\mu\sigma}I_{\alpha\beta\sigma}{}^{\nu}-\frac{1}{2}\eta^{\mu\nu}{\cal P}_{\alpha\beta\gamma\delta}\right)-\left(I_{\gamma\delta}{}^{\mu\nu}\eta_{\alpha\beta}l^{2}+I_{\alpha\beta}{}^{\mu\nu}\eta_{\gamma\delta}(l-q)^{2}\right)\right\}\right],\end{split}$ (8)
where
$I_{\mu\nu\lambda\rho}=\frac{1}{2}\left(\eta_{\mu\lambda}\eta_{\nu\rho}+\eta_{\mu\rho}\eta_{\nu\lambda}\right)$
We will use the standard integral [50]
$I^{r}_{s}=\int\frac{d^{d}k}{(2\pi)^{d}}\frac{(k^{2})^{r}}{(k^{2}-a^{2})^{s}}=\frac{i(-1)^{r+s}}{(4\pi)^{d/2}}\frac{1}{(a^{2})^{s-r-d/2}}\frac{\Gamma(r+d/2)\Gamma(s-r-d/2)}{\Gamma(d/2)\Gamma(s)},$ (9)
as well as the Fourier transforms
$\int\frac{d^{3}\vec{q}}{(2\pi)^{3}}\frac{e^{i\vec{q}\cdot\vec{r}}}{\vec{q}^{2}}=\frac{1}{4\pi r},\qquad\int\frac{d^{3}\vec{q}}{(2\pi)^{3}}\frac{e^{i\vec{q}\cdot\vec{r}}}{|\vec{q}|}=\frac{1}{2\pi^{2}r^{2}},\qquad\int\frac{d^{3}\vec{q}}{(2\pi)^{3}}e^{i\vec{q}\cdot\vec{r}}\ln\vec{q}^{2}=-\frac{1}{2\pi r^{3}}$ (10)
in the following calculations. Also, we will take the dummy virtual momentum named $l^{\mu}$ appearing below to run from -$\infty$ to $\infty$. The amplitudes will be evaluated in the usual relativistic framework, whereas the non-relativistic limit (such as $k_{1}\cdot k_{2}\approx-m_{1}m_{2}$) will be taken after the evaluation.